# Hessian Aware Low-Rank Weight Perturbation for Continual Learning

### v2

@@ Abstract
Continual learning aims to learn a series of tasks sequentially without forgetting the knowledge acquired from the previous ones. In this work, we propose the Hessian Aware Low-Rank Perturbation algorithm for continual learning. By modeling the parameter transitions along the sequential tasks with the weight matrix transformation, we propose to apply the low-rank approximation on the task-adaptive parameters in each layer of the neural networks. Specifically, we theoretically demonstrate the quantitative relationship between the Hessian and the proposed low-rank approximation. The approximation ranks are then globally determined according to the marginal increment of the empirical loss estimated by the layer-specific gradient and low-rank approximation error. Furthermore, we control the model capacity by pruning less important parameters to diminish the parameter growth. We conduct extensive experiments on various benchmarks, including a dataset with large-scale tasks, and compare our method against some recent state-of-the-art methods to demonstrate the effectiveness and scalability of our proposed method. Empirical results show that our method performs better on different benchmarks, especially in achieving task order robustness and handling the forgetting issue. The source code is at https://github.com/lijiaqi/HALRP.
@@
publicationid: pubid: $0$–$00$ 2023 IEEE
@@ 1 Introduction
To address this problem, some methods choose to expand the model during the dynamic learning process in a specific way, leading to inferencing with task-specific parameters for each task. For example, a recent work [9] proposed to decompose the model as task-private and task-shared parameters via an additive model parameters decomposition. However, this method only applies an attention vector for the masking of the task-shared parameters. Furthermore, the private parameters were controlled by a regularization and consolidation process, which lead to the linear model increment along with the increasing task number, damaging the scalability for deploying a CL system. Another solution to reduce the parameter increase is applying factorization for the model parameters. In this regard, [10] proposed a low-rank factorization method for the model parameters decomposition with a Bayesian process inference. However, this approach required a large rank for achieving desirable accuracy and also suffered from ineffectiveness in some complex data scenarios.
In this work, we propose a low-rank perturbation method to learn the relationship between the learned model (base parameters) and the parameters for new tasks. By formulating a flexible weight transition process, the model parameters during the CL scenario are decomposed as task-shared ones and task-adaptive ones. The former can be adopted from the base task to new tasks. The latter conforms to a low-rank matrix, and the number of relevant parameters can be effectively reduced for every single layer in a neural network across the tasks. With a simple warm-up training strategy, the low-rank task-adaptive parameters can be efficiently initialized with singular value decomposition.
@@
  • We proposed a Hessian aware low-rank perturbations framework that allows efficient memory and computation requests for learning continual tasks.
@@
  • Extensive experiments on several benchmarks and recent state-of-the-art baselines (e.g., regularization-, expansion-, and memory-based) were conducted to demonstrate the effectiveness of our method. The results testify the superiority of our HALRP, in terms of accuracy, computational efficiency, scalability, and task-order robustness.
@@ 2 Related Work
Regularization-based approaches diminish catastrophic forgetting by penalizing the parameter drift from the previous tasks using different regularizers [4, 5, 6]. Batch Ensemble [11] designed an ensemble weight generation method by the Hadamard product between a shared weight among all ensemble members and an ensemble member-specific rank-one matrix. In [12], the regularizer on the adaptive group sparsity was used to guarantee the exact sparsity and freezing of the model. [13] indicated that learning tasks in different low-rank vector subspaces orthogonal to each other can minimize task interferences. Functional regularizer [6], a Bayesian inference framework over the function space, tried to construct and memorize an approximate posterior belief over the underlying task-specific function. [14] applied a regularizer with decoupled prototype-based loss, which can improve the intra-class and inter-class structure significantly. Compared to this kind of approach, our method applied task-specific parameters and introduced an explicit weight transition process to leverage the knowledge from the previous tasks and overcome the forgetting issue.
@@
Expansion-based methods utilize different subsets of model parameters for each task. [15] proposed to memorize the learned knowledge by freezing the base model and progressively expanding the new sub-model for new tasks. In [16], the reuse and expansion of networks were achieved dynamically by selective retraining, with splitting or duplicating components for newly coming tasks. Some work (e.g.[17]) tried to determine the optimal network growth by neural architecture search. To balance memory stability and learning plasticity, [18] adopted a distillation-based method under the class incremental scenario. To achieve the scalability and the robustness of task orders, Additive Parameter Decomposition (APD) [9] adopted sparse task-specific parameters for novel tasks in addition to the dense task-shared ones and performed hierarchical consolidation within similar task groups for the further knowledge sharing. More recently, Winning Subnetworks (WSN) [19] jointly learned shared network weights and binary masks for each task to avoid retraining. To address the forgetting issue, only the model weights that had not been selected in the previous tasks were tendentiously updated during training. However, this method still needs to store task-specific mask information for the inference stage and relies on extra compression processes for mask encoding to achieve scalability.
@@
Replay (memory) based approaches usually leverage different types of replay buffers (e.g., [20, 21, 13, 22, 23, 24, 25]) to memorize a small episode of the previous tasks, which can be rehearsed when learning the novel ones to avoid forgetting. In [25], a dynamic prototype-guided memory replay module was incorporated with an online meta-learning framework to reduce memory occupation. [26] proposed to process both specific and generalized information by the interplay of three memory networks. To avoid the repeated inferences on previous tasks, Gradient Episodic Memory [2] and its variant [27] projected the new gradients into a feasible region that is determined by the gradients on previous task samples. Instead, Gradient Projection Memory (GPM) [2] chose to directly store the bases of previous gradient spaces to guide the direction of parameters update on new tasks. Compared to these approaches, our proposed method does not rely on extra storage for the data or gradient information of previous tasks, avoiding privacy leakage under some sensitive scenarios.
@@
Low-rank Factorization for CL Low-rank factorization [28] has been widely studied in deep learning to decompose the parameters for model compression [29, 30] or data projection [31, 32]. In the context of CL, [10] considers the low-rank model factorization and automatic rank selection per task for variational inference, which requires significantly large rank increments per task to achieve high accuracy. GPM  [2] applied the singular value decomposition on the representations and store them in the memory. [13] proposed to learn tasks by low-rank vector sub-spaces to avoid a joint vector space that may lead to interferences among tasks. The most similar work is Incremental Rank Updates (IRU) [33]. Compared to the decomposition in IRU, we adopted the low-rank approximation on the residual representation in the weights transitions. Furthermore, the rank selection in our work was dynamically and automatically determined according to Hessian-aware perturbations, rather than the manual rank increment in [33].
@@
$\mathrm{diag}(\boldsymbol{\Sigma})$ diagonal elements of $\boldsymbol{\Sigma}$
$\mathrm{Diag}(\sigma_{1},\cdots,\sigma_{n})$ diagonal matrix with elements $\{\sigma_{1},\cdots,\sigma_{n}\}$
@@ 3 Preliminary and Background
This section introduces the basic knowledge of continual learning and singular value decomposition [34]. In Table I, we provide a list of the main notations used in our paper. Specifically, the bold uppercase (or lowercase) letters indicate the matrices (or vectors) in the remaining part.
@@ 3.2 Low-Rank Approximation of Matrix with SVD
The singular value decomposition (SVD) factorizes a rectangular matrix $\mathbf{B}\in\mathbb{R}^{J\times I}$ with three matrices [^1] In this paper, we assume the eigenvalues (or the diagonal entries of $\boldsymbol{\Sigma}$) are always sorted with descending order in such an SVD decomposition, i.e., $\sigma_{1}\geq\sigma_{2}\geq...\geq\sigma_{r}$.: $\mathbf{B}=\mathbf{U}\boldsymbol{\Sigma}\mathbf{V}^{\top}$, where $\mathbf{U}\in\mathbb{R}^{J\times J}$, $\mathbf{V}\in\mathbb{R}^{I\times I}$, and $\boldsymbol{\Sigma}\in\mathbb{R}^{J\times I}$. Denote the rank of $\mathbf{B}$ with $r\leq\min\{I,J\}$, then $\mathbf{B}$ can be further expressed as $\mathbf{B}=\sum_{i=1}^{r}\sigma_{i}\mathbf{u}_{i}\mathbf{v}_{i}^{\top}$, where $\sigma_{i}\in\mathrm{diag}(\boldsymbol{\Sigma})$ is the singular values, $\mathbf{u}_{i}$ and $\mathbf{v}_{i}$ are respectively the left and right singular vectors. In this work, we take the property of $k$-rank approximation of $\mathbf{B}$ (with $k\leq r$) by leveraging the Eckart–Young–Mirsky theorem [35], which can be expressed with top-$k$ leading singular values and vectors: $\mathbf{B}^{(k)}=\mathbf{U}^{(k)}\boldsymbol{\Sigma}^{(k)}\mathbf{V}^{(k)\top}=\sum_{i=1}^{k}\sigma_{i}\mathbf{u}_{i}\mathbf{v}_{i}^{\top}$, where $\mathbf{U}^{(k)}$, $\boldsymbol{\Sigma}^{(k)}$, $\mathbf{V}^{(k)}$ are the corresponding leading principal sub-matrices [^2] With a slight abuse of notation, the notation $\mathbf{M}^{(k)}$ refers to the $k$-rank approximation of the matrix $\mathbf{M}$ according to the context: (1) $\mathbf{U}^{(k)},\boldsymbol{\Sigma}^{(k)},\mathbf{V}^{(k)}$ indicate the top-$k$ leading submatrices; (2) $\mathbf{B}^{(k)}$ means the $r$-rank low-rank approximation for $\mathbf{B}$, then same for $\mathbf{W}^{(k)}$.. $k$ is chosen to tolerate a certain approximation error under the Frobenius norm $\|\cdot\|_{F}$ (See Appendix A for the proof):
$\begin{split}\|\mathbf{B}-\mathbf{B}^{(k)}\|_{F}=\sqrt{\sigma_{k+1}^{2}+\dots+\sigma_{r}^{2}}\end{split}$ (1)
@@ 4.1 Fully Connected Layers
We first consider linear layers of neural networks in a supervised training context. We begin by learning task $\mathcal{T}_{0}$ without any constraints on the model parameters. Specifically, we train the model by minimizing the empirical risk to get the base weights $\mathbf{W}^{\text{base}}=\operatorname*{arg\,min}_{\mathbf{W}}\mathcal{L}(\mathbf{W};\mathcal{D}_{0})$, where $\mathbf{W}^{\text{base}}\in\mathbb{R}^{J\times I}$, $J$ is the output dimension, and $I$ is the input dimension of the layer.
@@
where $\mathbf{B}_{1}\in\mathbb{R}^{J\times I}$ is a sparse low-rank matrix, $\mathbf{R}_{1}=\textrm{Diag}(r_{1,0},\dots,r_{1,J})$, and $\mathbf{S}_{1}=\textrm{Diag}(s_{1,0},\dots,s_{1,I})$ are task-private parameters for $\mathcal{T}_{1}$ [^3] We will show later that $\mathbf{B}$ can be further approximated by $\mathbf{U}^{(k)},\boldsymbol{\Sigma}^{(k)},\mathbf{V}^{(k)}$ with low-rank approximation, so we can write the task-private weights for task $t$ as either $\{\mathbf{R}_{t},\mathbf{S}_{t},\mathbf{B}_{t}^{(k)}\}$ or $\{\mathbf{R}_{t},\mathbf{S}_{t},\mathbf{U}_{t}^{(k)},\boldsymbol{\Sigma}_{t}^{(k)},\mathbf{V}_{t}^{(k)}\}$..
@@
In addition, regarding the parameter estimation, we can substitute $\mathbf{B}_{1}$ with its $k$-rank decomposition $\mathbf{B}_{1}^{(k)}=\sum_{i=1}^{k}\sigma_{i}\mathbf{u}_{i}\mathbf{v}_{i}^{\top}$ in Eq. 2 and directly minimize the empirical risk on task $\mathcal{T}_{1}$ to get the parameter estimation through stochastic gradient descent with the random initialization of $\{\mathbf{u}_{i}\}_{i=1}^{k}$ and $\{\mathbf{v}_{i}\}_{i=1}^{k}$. However, in practice, we found it challenging to learn a useful decomposition in this way (i.e., the estimations converge to a local optimum quickly) and the empirical risk minimization does not benefit from the low-rank decomposition. Due to the homogeneity among the one-rank components $\{\mathbf{u}_{i}\mathbf{v}_{i}^{\top}\}_{i=1}^{k}$, random initialization cannot sufficiently distinguish them and hence make the gradient descent ineffective.
@@
Eq. 4 can be solved efficiently via LSE minimization. For example, for $\mathbf{R}$, we have $\|\mathbf{W}^{\text{free}}-\mathbf{R}\mathbf{W}^{\text{base}}\|_{F}^{2}=\sum_{i=1}^{I}\sum_{j=1}^{J}(w^{\text{free}}_{ji}-r_{j}w^{\text{base}}_{ji})^{2}$, where $w^{\text{free}}_{ji}$ and $w^{\text{base}}_{ji}$ are the elements of $j$-th row and $i$-th column of $\mathbf{W}^{\text{free}}$ and $\mathbf{W}^{\text{base}}$, respectively. Applying the derivative w.r.t. $r_{j}$ and let it equals 0, we have $r_{j}=\sum_{i}w_{ji}^{\text{free}}w_{ji}^{\text{base}}/\sum_{i}(w_{ji}^{\text{base}})^{2}$. Then, we can solve $\mathbf{S}$ and $\mathbf{B}$ (see Appendix C for more details). Furthermore, with the SVD solver denoted by $\texttt{SVD}(\cdot)$, the values of the low-rank decomposition for $\mathbf{B}^{\text{free}}$ can be obtained:
$\begin{split}\mathbf{U}_{1}^{\text{free}},\boldsymbol{\Sigma}_{1}^{\text{free}},\mathbf{V}_{1}^{\text{free}}&=\texttt{SVD}(\mathbf{B}_{1}^{\text{free}})\end{split}$ (5)

### v5

# Hessian Aware Low-Rank Perturbation for Order-Robust Continual Learning

@@ Abstract
Continual learning aims to learn a series of tasks sequentially without forgetting the knowledge acquired from the previous ones. In this work, we propose the Hessian Aware Low-Rank Perturbation algorithm for continual learning. By modeling the parameter transitions along the sequential tasks with the weight matrix transformation, we propose to apply the low-rank approximation on the task-adaptive parameters in each layer of the neural networks. Specifically, we theoretically demonstrate the quantitative relationship between the Hessian and the proposed low-rank approximation. The approximation ranks are then globally determined according to the marginal change of the empirical loss estimated by the layer-specific gradient and low-rank approximation error. Furthermore, we control the model capacity by pruning less important parameters to diminish the parameter growth. We conduct extensive experiments on various benchmarks, including a dataset with large-scale tasks, and compare our method against some recent state-of-the-art methods to demonstrate the effectiveness and scalability of our proposed method. Empirical results show that our method performs better on different benchmarks, especially in achieving task order robustness and handling the forgetting issue. The source code is at https://github.com/lijiaqi/HALRP.
@@
publicationid: pubid: This work has been submitted to the IEEE for possible publication. Copyright may be transferred without notice, after which this version may no longer be accessible.
@@ 1 Introduction
To address this problem, some methods choose to expand the model during the dynamic learning process in a specific way, leading to inferencing with task-specific parameters for each task. For example, a recent work [9] proposed to decompose the model as task-private and task-shared parameters via an additive model parameters decomposition. However, this method only applies an attention vector for the masking of the task-shared parameters. Furthermore, the private parameters were controlled by a regularization and consolidation process, which lead to the linear model increment along with the increasing task number, damaging the scalability for deploying a CL system. Another solution to reduce the parameter increase is to apply factorization for the model parameters. In this regard, [10] proposed a low-rank factorization method for the model parameters decomposition with a Bayesian process inference. However, this approach required a large rank for achieving desirable accuracy and also suffered from ineffectiveness in some complex data scenarios.
In this work, we propose a low-rank perturbation method to learn the relationship between the learned model (base parameters) and the parameters for new tasks. By formulating a flexible weight transition process, the model parameters during the CL scenario are decomposed as task-shared ones and task-adaptive ones. The former can be adopted from the base task to new tasks. The latter conforms to a low-rank matrix, and the number of relevant parameters can be effectively reduced for every single layer in a neural network across the tasks. With a simple warm-up training strategy, low-rank task adaptive parameters can be efficiently initialized with singular value decomposition.
@@
  • We proposed a Hessian-aware low-rank perturbations framework that allows efficient memory and computation requests for learning continual tasks.
@@
  • Extensive experiments on several benchmarks and recent state-of-the-art baselines (e.g., regularization-, expansion-, and replay-based) were conducted to demonstrate the effectiveness of our method. The results testify the superiority of our HALRP, in terms of accuracy, computational efficiency, scalability, and task-order robustness.
@@ 2 Related Work
Regularization-based approaches diminish catastrophic forgetting by penalizing the parameter drift from the previous tasks using different regularizers [4, 5, 6]. Batch Ensemble [11] designed an ensemble weight generation method by the Hadamard product between a shared weight among all ensemble members and an ensemble member-specific rank-one matrix. [12] indicated that learning tasks in different low-rank vector subspaces orthogonal to each other can minimize task interferences. [13] applied a regularizer with decoupled prototype-based loss, which can improve the intra-class and inter-class structure significantly. Compared to this kind of approach, our method applied task-specific parameters and introduced an explicit weight transition process to leverage the knowledge from the previous tasks and overcome the forgetting issue.
@@
Expansion-based methods utilize different subsets of model parameters for each task. [14] proposed to memorize the learned knowledge by freezing the base model and progressively expanding the new sub-model for new tasks. In [15], the reuse and expansion of networks were achieved dynamically by selective retraining, with splitting or duplicating components for newly coming tasks. Some work (e.g.[16]) tried to determine the optimal network growth by neural architecture search. To balance memory stability and learning plasticity, [17] adopted a distillation-based method under the class incremental scenario. To achieve the scalability and the robustness of task orders, Additive Parameter Decomposition (APD) [9] adopted sparse task-specific parameters for novel tasks in addition to the dense task-shared ones and performed hierarchical consolidation within similar task groups for the further knowledge sharing.  [18] introduced Channel-Wise Linear Reprogramming (CLR) transformations on the output of each convolutional layer of the base model as the task-private parameters. However, this method relied on the prior knowledge from a disjoint dataset (e.g., ImageNet-1K). Winning Subnetworks (WSN) [19] jointly learned the shared network and binary masks for each task. To address the forgetting issue, only the model weights that had not been selected in the previous tasks were tendentiously updated during training. But this method still needs to store task-specific masks for the inference stage and relies on extra compression processes for mask encoding to achieve scalability.
@@
Replay-/Memory-based approaches usually leverage different types of replay buffers (e.g., [20, 21, 12, 22, 23, 24, 25]) to memorize a small episode of the previous tasks, which can be rehearsed when learning the novel ones to avoid forgetting. In [25], a dynamic prototype-guided memory replay module was incorporated with an online meta-learning framework to reduce memory occupation. [26] proposed to process both specific and generalized information by the interplay of three memory networks. To avoid the repeated inferences on previous tasks, Gradient Episodic Memory [27] and its variant [28] projected the new gradients into a feasible region that is determined by the gradients on previous task samples. Instead, Gradient Projection Memory (GPM) [2] chose to directly store the bases of previous gradient spaces to guide the direction of parameters update on new tasks. Compared to these approaches, our proposed method does not rely on extra storage for the data or gradient information of previous tasks, avoiding privacy leakage under some sensitive scenarios.
@@
Low-rank Factorization for CL Low-rank factorization [29] has been widely studied in deep learning to decompose the parameters for model compression [30, 31] or data projection [32, 33]. In the context of CL, [10] considers the low-rank model factorization and automatic rank selection per task for variational inference, which requires significant large rank increments per task to achieve high accuracy. GPM [2] applied the singular value decomposition on the representations and store them in the memory. [12] proposed to learn tasks by low-rank vector sub-spaces to avoid a joint vector space that may lead to interferences among tasks. The most similar work is Incremental Rank Updates (IRU) [34]. Compared to the decomposition in IRU, we adopted the low-rank approximation on the residual representation in the weights transitions. Furthermore, the rank selection in our work was dynamically and automatically determined according to Hessian-aware perturbations, rather than the manual rank increment in [34].
@@
$\mathrm{diag}(\boldsymbol{\Sigma})$ the set of diagonal elements of $\boldsymbol{\Sigma}$
$\mathrm{Diag}(\sigma_{1},\cdots,\sigma_{n})$ diagonal matrix elements $\{\sigma_{1},\cdots,\sigma_{n}\}$
@@ 3 Preliminary and Background
This section introduces the basic knowledge of continual learning and singular value decomposition. In Table I, we provide a list of the main notations used in our paper. Specifically, the bold uppercase (or lowercase) letters indicate the matrices (or vectors) in the remaining part.
@@ 3.2 Low-Rank Approximation of Matrix with SVD
The singular value decomposition (SVD) factorizes a rectangular matrix $\mathbf{B}\in\mathbb{R}^{J\times I}$ with three matrices [^1] In this paper, we assume the eigenvalues are always sorted with descending order in such SVD decompositions, i.e., $\sigma_{1}\geq\sigma_{2}\geq...\geq\sigma_{r}$.: $\mathbf{B}=\mathbf{U}\boldsymbol{\Sigma}\mathbf{V}^{\top}$, where $\mathbf{U}\in\mathbb{R}^{J\times J}$, $\mathbf{V}\in\mathbb{R}^{I\times I}$, and $\boldsymbol{\Sigma}\in\mathbb{R}^{J\times I}$. Denote the rank of $\mathbf{B}$ with $r\leq\min\{I,J\}$, then $\mathbf{B}$ can be further expressed as $\mathbf{B}=\sum_{i=1}^{r}\sigma_{i}\mathbf{u}_{i}\mathbf{v}_{i}^{\top}$, where $\sigma_{i}\in\mathrm{diag}(\boldsymbol{\Sigma})$ is the singular values, $\mathbf{u}_{i}$ and $\mathbf{v}_{i}$ are respectively the left and right singular vectors. In this work, we take the property of $k$-rank approximation of $\mathbf{B}$ (with $k\leq r$) by leveraging the Eckart–Young–Mirsky theorem [35], which can be expressed with top-$k$ leading singular values and vectors: $\mathbf{B}^{(k)}=\mathbf{U}^{(k)}\boldsymbol{\Sigma}^{(k)}(\mathbf{V}^{(k)})^{\top}=\sum_{i=1}^{k}\sigma_{i}\mathbf{u}_{i}\mathbf{v}_{i}^{\top}$, where $\mathbf{U}^{(k)}$, $\boldsymbol{\Sigma}^{(k)}$, $\mathbf{V}^{(k)}$ are the corresponding leading principal sub-matrices [^2] With a slight abuse of notation, the notation $\mathbf{M}^{(k)}$ refers to the $k$-rank approximation of the matrix $\mathbf{M}$ according to the context: (1) $\mathbf{U}^{(k)},\boldsymbol{\Sigma}^{(k)},\mathbf{V}^{(k)}$ indicate the top-$k$ leading submatrices; (2) $\mathbf{B}^{(k)}$ means the $r$-rank low-rank approximation for $\mathbf{B}$, then same for $\mathbf{W}^{(k)}$.. $k$ is chosen to tolerate a certain approximation error under the Frobenius norm $\|\cdot\|_{F}$ (See Appendix A for the proof):
$\begin{split}\|\mathbf{B}-\mathbf{B}^{(k)}\|_{F}=\sqrt{\sigma_{k+1}^{2}+\dots+\sigma_{r}^{2}}.\end{split}$ (1)
@@ 4.1 Fully Connected Layers
We first consider linear layers of neural networks. We begin by learning task $\mathcal{T}_{0}$ without any constraints on the model parameters. Specifically, we train the model by minimizing the empirical risk to get the base weights $\mathbf{W}^{\text{base}}=\operatorname*{arg\,min}_{\mathbf{W}}\mathcal{L}(\mathbf{W};\mathcal{D}_{0})$, where $\mathbf{W}^{\text{base}}\in\mathbb{R}^{J\times I}$, $J$ is the output dimension, and $I$ is the input dimension of the layer.
@@
where $\mathbf{B}_{1}\in\mathbb{R}^{J\times I}$ is a sparse low-rank matrix, $\mathbf{R}_{1}=\textrm{Diag}(r_{1},\dots,r_{J})$, and $\mathbf{S}_{1}=\textrm{Diag}(s_{1},\dots,s_{I})$ are task-private parameters for $\mathcal{T}_{1}$ [^3] We will show later that $\mathbf{B}$ can be further approximated by $\mathbf{U}^{(k)},\boldsymbol{\Sigma}^{(k)},\mathbf{V}^{(k)}$ with low-rank approximation, so we can write the task-private weights for task $t$ as either $\{\mathbf{R}_{t},\mathbf{S}_{t},\mathbf{B}_{t}^{(k)}\}$ or $\{\mathbf{R}_{t},\mathbf{S}_{t},\mathbf{U}_{t}^{(k)},\boldsymbol{\Sigma}_{t}^{(k)},\mathbf{V}_{t}^{(k)}\}$..
@@
In addition, regarding the parameter estimation, we can substitute $\mathbf{B}_{1}$ with its $k$-rank approximation $\mathbf{B}_{1}^{(k)}=\sum_{i=1}^{k}\sigma_{i}\mathbf{u}_{i}\mathbf{v}_{i}^{\top}$ in Eq. 2 and directly minimize the empirical risk on task $\mathcal{T}_{1}$ to get the parameter estimation through stochastic gradient descent with the random initialization of $\{\mathbf{u}_{i}\}_{i=1}^{k}$ and $\{\mathbf{v}_{i}\}_{i=1}^{k}$. However, in practice, we found it challenging to learn a useful decomposition in this way (i.e., the estimations converge to a local optimum quickly) and the empirical risk minimization does not benefit from the low-rank decomposition. Due to the homogeneity among the one-rank components $\{\mathbf{u}_{i}\mathbf{v}_{i}^{\top}\}_{i=1}^{k}$, random initialization cannot sufficiently distinguish them and hence make the gradient descent ineffective.
@@
Eq. 4 can be solved efficiently via LSE minimization. For example, for $\mathbf{R}$, we have $\|\mathbf{W}^{\text{free}}-\mathbf{R}\mathbf{W}^{\text{base}}\|_{F}^{2}=\sum_{i=1}^{I}\sum_{j=1}^{J}(w^{\text{free}}_{ji}-r_{j}w^{\text{base}}_{ji})^{2}$, where $w^{\text{free}}_{ji}$ and $w^{\text{base}}_{ji}$ are the elements of $j$-th row and $i$-th column of $\mathbf{W}^{\text{free}}$ and $\mathbf{W}^{\text{base}}$, respectively. Applying the derivative w.r.t. $r_{j}$ and let it equal 0, we have $r_{j}=(\sum_{i}w_{ji}^{\text{free}}w_{ji}^{\text{base}})/\sum_{i}(w_{ji}^{\text{base}})^{2}$. Then, we can solve $\mathbf{S}$ and $\mathbf{B}$ (see Appendix C for more details). Furthermore, with the SVD solver denoted by $\texttt{SVD}(\cdot)$, the values of the low-rank decomposition for $\mathbf{B}^{\text{free}}$ can be obtained:
$\begin{split}\mathbf{U}_{1}^{\text{free}},\boldsymbol{\Sigma}_{1}^{\text{free}},\mathbf{V}_{1}^{\text{free}}&\leftarrow\texttt{SVD}(\mathbf{B}_{1}^{\text{free}})\end{split}$ (5)